Find the following indefinite integrals: x+1 dx (a) S *7+7x (b) S(Vx? – 5) dx solving step by step
Answer #1

f x6+1 dx 27+70 take x7+7x=t (726+7) dx = at 7 (2671) dx=dt (6+1)dx= at 7 SO: Now inegral will be: at 7 t f at t ent to ㅋ putHora - 54)dx f (x = 3 -5%) dx a Apply linearity: f 4 dx 15 5* da 7/t 24 - 154 dx It dn = - ak so he get apply faxar en a x xfeel free to ask if you have any doubt and if you liked my answer please give a thumbs up